OpenStudy (anonymous):

factor completely(14x^-5)-(35x^-6)

OpenStudy (anonymous):

I just need someone to let me know if this answer is correct (2x-5)/(7x^6)

OpenStudy (lalaly):

its 7(2x-5) /x^6

OpenStudy (anonymous):

why not 7x^6 on the bottom

hero (hero):

\[\frac{14}{x^5} - \frac{35}{x^6}\]

hero (hero):

is that the original problem?

OpenStudy (lalaly):

because only the x has the power -6

OpenStudy (anonymous):

but i pulled 7x^-6 out as a whole

hero (hero):

\[\frac{7}{x^5}(2 - \frac{5}{x})\]

hero (hero):

That's what I get